find the slope for (7,-2)&(-5,1)
Answer by sdmmadam@yahoo.com(530)

find the slope for (7,-2)&(-5,1)
find the slope for (7,-2)&(-5,1)
The slope of the line joining two given points P(x1,y1) and Q(x2,y2)
is given by
slope = m = (y2-y1)/(x2-x1) ----(*)
Therefore the slope of the line joining P(7,-2) and Q(-5,1)
is given by
slope = [1-(-2)]/[(-5)-(7)]= (1+2)/(-5-7) = 3/(-12) = (-1/4)
Answer: slope = (-1/4)
Note: Here P(x1,y1)=P(7,-2) giving x1=7 and y1= -2
and Q(x2,y2)=Q(-5,1) giving x2=-5 and y2=1